### v2

# Pressure Induced Altermagnetism in Layered Ternary Iron-Selenides

Abstract
Employing first-principles based calculations, we reexamined the high-pressure phases of the vacancy-ordered iron-selenides, i.e. Fe4Se5 phase. A magnetic transition from the block-spin antiferromagnetic phase to Néel-AM phase is observed under high pressure when the iron-vacancy order is preserved. The transition is first-order, driven by the collapse of -axis and accompanied by an insulator-metal transition. In addition, the Néel-AM phase is a rare example of intrinsic altermagnetism that does not require ligand atoms nor supercell construction to break the spatial inversion or translational symmetry between the two spin sublattices, with a spin-splitting of the band structure as large as 300 meV. If the re-entrant superconducting phase of KyFe2-xSe2 under high pressure emerges from the Néel-AM normal state, an equal-spin triplet pairing would be naturally favored.
Introduction. The concept of altermagnetism has attracted considerable attention recentlyŠmejkal et al. (2022a). Theoretically, altermagnetism is closely related to and resembles anisotropic Fermi-liquid instabilitiesWu and Zhang (2004); Wu et al. (2007), and is characterized by spin-orbit coupling (SOC) free compensated spin-up and spin-down sublattices that are spatially connected by rotational symmetryŠmejkal et al. (2022b); Chen et al. (2024), instead of the translational or inversion symmetry in conventional antiferromagnetism (AFM). Therefore, the spin-degeneracy is lifted in the altermagnetism at general -points, since the or symmetry is brokenŠmejkal et al. (2022b); Chen et al. (2024); Li et al. (2022); Krempaský et al. (2024) (, and represent time-reversal, lattice translation and inversion symmetry, respectively). Such feature is desirable in spintronicsChen et al. (2022); Šmejkal et al. (2022a), and may lead to exotic superconducting pairing statesMazin (2022); Šmejkal et al. (2022a); Zhang et al. (2024); Mæland et al. (2024); Wei et al. (2024); Chakraborty and Black-Schaffer (2024); Sukhachov et al. (2025). However, material candidates for altermagnetism-related superconductivity (SC) is currently still lacking and remain elusive.
Of the iron-based superconductors, the ternary iron-selenide compounds Fe2-xSe2 (=Tl, K, Rb, Cs) are of particular interestGuo et al. (2010); Wang et al. (2011); Fang et al. (2011). Structurally, these compounds are similar to the 122-type iron-pnictidesSasmal et al. (2008). The anti-PbO-type FeSe layers are intercalated by atoms, leaving certain amount of Fe-vacancies in the Fe-square lattice(FIG. 1). At , a robust vacancy-ordered superstructure forms, and the ground state becomes block-spin AFM (BS-AFM)Bao et al. (2011); Cao and Dai (2011); Yan et al. (2011); Ye et al. (2011) (FIG. 1c). In addition, the compound is insulating at Cao and Dai (2011); Yan et al. (2011). Therefore, the Fe4Se5(Fe1.6Se2) distinguishes itself from rest of the family and is often referred to as the 245-phase. The ordered Fe-vacancy is found to be quite robust and universally exists not only in Fe2-xSe2 and Fe2-xS2Toulemonde et al. (2013); Bao (2014); Mangelis et al. (2019); Croitori et al. (2020); Chen et al. (2021), but also in -Fe1-xSeChen et al. (2014); Gao et al. (2017); Yeh et al. (2020), even under very high pressure up to 16 GPaSun et al. (2012); Ksenofontov et al. (2012). In addition, a re-emerging superconductiviting phase between 10.5 GPa and 13.2 GPa was observed experimentallySun et al. (2012). However, the superconducting phase of Fe2-xSe2 was the focus of previous studies, and much less attention was paid to the insulating 245-phase. The stability of such vacancy order and its influence on the magnetic as well as the insulator-metal transitions in this material family deserve more investigations.
In this paper, we reexamine the high-pressure phase of the vacancy-ordered 245-phase selenides, i.e. Fe4Se5. We found a first-order transition from the BS-AFM phase to the Néel-AM phase in all these compounds in addition to (Tl, K) we found earlierCao et al. (2011), under certain hydraulitic pressure . We address the issue of overestimated magnetic interactions using DFT+DMFT calculationsKotliar et al. (2006); Haule (2007), and found the competing stripe-like AFM phase is not favored. The magnetic phase transition is driven by the collapse of -axis, and is accompanied by an insulator-metal transition. In addition, the Néel-AM phase is altermagnetic, exhibiting spin-splitted band structure up to 300 meV. Unlike RuO2Šmejkal et al. (2020); Shao et al. (2021) or other proposed altermagnetic compoundsŠmejkal et al. (2022b), the Néel-AM phase does not require ligand atoms to break the inversion or translation symmetry between the spin-sublattices, and its magnetic unit cell is the same size as its nonmagnetic primitive cell. As a result, the spin-splitting in the Néel-AM phase is more robust since it does not depend on the ligand atom positions that are susceptible to defect, external pressure, temperature, or doping. Therefore, the Néel-AM pattern is an example of intrinsic altermagnetism. This renders the high-pressure superconducting phase of KyFe2-xSe2 between 10.5 GPa and 13.2 GPa a possible candidate for studying inter-play between superconductivity and altermagnets.
Magnetic Phase Transition under Pressure. Employing a high-throughput methodXu et al. (2025), we have searched the ground state magnetic phase of Fe4Se5 under different pressure. There are 8 inequivalent AFM configurations for the primitive cell (8 Fe atoms) (see SI section B and FIG. S-2 for complete list) and 508 inequivalent AFM configurations for size-2 supercell (16 Fe atoms). Previously, an extended model has been identified to well explain the magnetic configurationsCao and Dai (2011); Ke et al. (2012); Chi et al. (2013), suggesting that the magnetic interactions are fairly local in the 245 phase. In this model, the effective spin Hamiltonian can be written as: , where and denote intra-block and inter-block nearest neighbors, and are intra-block and inter-block next nearest neighbors, respectively. Therefore, we constrain the search within the primitive cell in most cases. In addition, neutron scattering experiment also reveals small inter-layer interactions (1.0 meV)Chi et al. (2013), thus we consider in-plane configurations only. The stability of the Néel-AM ground state is further confirmed based on extended parameters at 12 GPa using Monte Carlo simulations(see SI section C and FIG. S-3 for details).
The BS-AFM phase is the ground state for all the compounds under ambient pressure, and is energetically overwhelmingly favored. For example, the second lowest energy phase is 280 meV/f.u. higher than the BS-AFM phase for Rb2Fe4Se5, and the Néel-AM phase is 340 meV/f.u. higher than the BS-AFM phase (FIG. 2a). At 10 GPa, the BS-AFM remains as the ground state, while the Néel-AM phase becomes second lowest, and its enthalpy is only 17 meV/f.u. higher. At 11 GPa and beyond (to at least 16 GPa), the Néel-AM phase becomes the lowest enthalpy state. Between 12 GPa and 16 GPa, its enthalpy is at least 30 meV/f.u. lower than the second lowest state. While a stripe-like AFM was shown to be a low-energy phase and compete with the Néel-AM phase at the DFT levelCao et al. (2011), the magnetism of the stripe-like AFM phase is known to be severely overestimated in normal DFT calculations, and the calculated staggered moment is several times larger than experimental valueMazin et al. (2008). This is in sharp contrast to the BS-AFM phase whose magnetic moments from DFT calculations are in good agreement with experimentBao et al. (2011); Cao and Dai (2011). In addition, our DFT+DMFT calculations at 12 GPa show that the enthalpy of the stripe-like AFM phase is in fact 100 meV/f.u. higher than the Néel-AM phase, and the ordered moments are consistent with experimental observation (see SI section E for details). We also confirm that no AFM configuration for size-2 supercell has lower enthalpy than Néel-AM phase at 12 GPa. Thus, we conclude that the ground state magnetic phase is Néel-AM between 11 and 16 GPa in Rb2Fe4Se5. Similarly, this phase transition can be identified around 11 GPa in K2Fe4Se5, 13 GPa in Na2Fe4Se5, 12 GPa in Tl2Fe4Se5, and 9 GPa in Cs2Fe4Se5, respectively (see SI section F and FIG S-7 for details).
External pressure effect is also evident on the magnetic moment. Under ambient pressure, Rb2Fe4Se5 exhibits 2.9 /Fe ordered moment (FIG. 2b), in good agreement with experimental observation of 3.31 /Fe Bao et al. (2011). It is almost linearly suppressed to 2.56 /Fe at 10 GPa before the transition, then abruptly reduced to 1.91 /Fe in the Néel-AM phase at 11 GPa. Further increasing the pressure, the ordered moment is again linearly suppressed to 1.55 /Fe at 16 GPa. We also note that the AFM-51, AFM-54, AFM-99, AFM-83 and AFM-86 phases become unstable between 8 and 14 GPa, and AFM-53 becomes unstable at 16 GPa, as their ordered moments are quickly suppressed to less than 0.1 /Fe and the total enthalpies approach to NM phases. Therefore, these phases become nonmagnetic at high pressure in calculation. The ordered moment is robust in both the BS-AFM and the Néel-AM phase, as no sign of sudden drop can be observed up to 16 GPa. The result also supports a competition only between the BS-AFM phase and the Néel-AM phase under high pressure.
Evolution of Crystal Structure. The space group of fully relaxed BS-AFM phase Rb2Fe4Se5 at ambient pressure is 4/m (No. 87). Each primitive cell contains 2 chemical formula units, and all 8 Fe-sites are equivalent. The lattice constants and of Rb2Fe4Se5 are 8.730Å and 14.888Å under ambient pressure in the BS-AFM phase, respectively. These values are in good agreement with the experimental observationYe et al. (2011), with less than 2% error. At 8 GPa, and are reduced by 5.0% and 8.9% compared to and , respectively, which is also in perfect agreement with the experimental observationYe et al. (2014). From 10 GPa to 11 GPa across the magnetic phase transition, the space group remains unchanged. However, the lattice constant is suddenly reduced by 4.4% while the change in is less than 0.3%. Thus, the structural transition is signaled by the collapse of -axis. The -collapse effect under pressure was indeed reported in KyFe2-xS2Tsuchiya et al. (2017), and can be inferred from the discontinuity in in KyFe2-xSe2Bendele et al. (2013). Although such transition was attributed to formation of disordered vacancy phase previouslyBendele et al. (2013); Yamamoto et al. (2016), the transition from BS-AFM to Néel-AM phase provides an alternative explanation to the observed effect. Since both the deviation from stoichemical composition and thermal fluctuations enhance disorder effectBao (2014), a careful characterization of stoichemical semiconducting sample under pressure may resolve this issue. Furthermore, the dynamical stability of the Néel-AM phase is also verified by calculating its phonon spectrum at 12 GPa (see SI section C and FIG. S-4 for details).
Spin Model Analysis. We briefly discuss the formation of the Néel-AM phase under the framework of the spin models. For the vacancy-ordered 245 phase, the simplest spin model would be model that considers nearest neighboring interactions only but differentiates the intra-block () and inter-block () interactions (FIG. 1). In such case, it is apparent that 4 possible magnetic phases exist depending on the signs of and : checker-board AFM ( and ), BS-AFM ( and ), Néel-AM ( and ), and FM ( and ). Thus, both BS-AFM and Néel-AM phase are dominating phases depending on the overall nature of intra-block and inter-block spin-interactions. If the next-nearest neighboring interactions are included, i.e. the extended model, the above statement remains valid as shown by Yu et al. using the classical Monte-Carlo simulationsYu et al. (2011). Using the classical expressions for different phasesYu et al. (2011); Cao et al. (2011), the Néel-AM phase is energetically favored over the BS-AFM, checkerboard AFM and the FM phase, if , , and , respectively. In addition, the energy of the Néel-AM phase is lower than the stripe-like AFM if .
We fitted the exchange interactions s of the extended - model under different pressure (see SI section G and TAB. S-II for details). At ambient pressure, the fitted meV/, meV/, meV/ and meV/. These are in reasonable agreement with previous experimentChi et al. (2013) (FIG. 2d), although both inter-block interactions and seem to be substantially overestimated. Under pressure, is nearly unchanged, while both and slightly enhances below 6 GPa and gradually decreases. In contrast, the pressure drastically alters , which changes its sign close to 9 GPa, where all interactions become antiferromagnetic. Nevertheless, even at 10 GPa, is too small compared to to drive the magnetic phase transition. Thus, we believe that the phase transition is first-order, induced by the sudden reduction of . To confirm this, we have also performed calculations with reduced with fixed at 10 GPa. Both and are drastically reduced if is under stress, while quickly further enhances. The condition is fulfilled when the reaches -7%. It is therefore reasonable to postulate that a uniaxial pressure in -direction is more effective to tune the exchange interactions and the Néel-AM phase may be achieved at lower pressure in such case.
Evolution of Electronic Structure. Under ambient pressure, Rb2Fe4Se5 is an insulator with an indirect gap meV, formed by the Fe-3 states at X (, 0, 0) and the Fe-3 state at (FIG. S-8a). Under pressure, the band gap is monotonically suppressed to 290 meV at 10 GPa (FIG. S-8b). In contrast, the Néel-AM phase is always metallic, therefore the magnetic phase transition is accompanied by a insulator-metal transition simultaneously.
As shown in Fig. 1(d), the spin-up and spin-down sublattices of the Néel-AM phase are spatially connected by the rotation, thus it is altermagnetic. Similar to the “failed AF-SOD” phase in CsCr3Sb5Xu et al. (2025), the Néel-AM pattern does not require ligand atoms to break the inversion or translation symmetry between these two sublattices. However, unlike the “failed AF-SOD” phase, the magnetic unit cell of the Néel-AM phase is the same size as the chemical unit cell. Therefore, the Néel-AM phase is also different from the supercell altermagnetsJaeschke-Ubiergo et al. (2024), and we denote it as intrinsic altermagnetism. As shown in FIG. 3(a), the spin degeneracy is lifed at general -points in Néel-AM phase Rb2Fe4Se5, although the spatial inversion symmetry is preserved in the compound. Therefore, . The spin-splitting close to at X (, 0, 0) or Y (0, , 0) can be estimated by the Fe-3 states to be 300 meV at 14 GPa. Finally, in the Néel-AM phase, the electronic states close to the Fermi level are dominated by Fe-3 and Fe-3 orbitals.
Discussions. Our prediction of the Néel-AM phase is highly dependent on the formation of the vacancy-order. However, controversial results regarding whether the vacancy-order can survive under high pressure have been reportedSvitlyk et al. (2011); Sun et al. (2012); Ksenofontov et al. (2012); Ye et al. (2014); Tsuchiya et al. (2017). We note that these experiments were performed with superconducting samples, whose composition deviates from the perfect 245-phase. In addition, the iron vacancy-ordered phase is much more favored at low temperaturesSvitlyk et al. (2011); Ye et al. (2014). Nevertheless, the synchrotron results of 245 selenides are in agreement that the structure persists to at least 12 GPaSvitlyk et al. (2011); Ksenofontov et al. (2012). Therefore, it is possible to realize the Néel-AM phase under sufficiently high pressure in 245 selenides. In fact, previous high pressure neutron scattering experiment has observed disappearance of the magnetic Bragg peak at lower pressure than the vacancy-order Bragg peakYe et al. (2014). We note here that the propagation wave vector of Néel-AM phase is (Cao et al., 2011), and always coincide with the structural peaks. Therefore, the disappearance of the magnetic peaks does not preclude the existence of the Néel-AM phase.
It is also worth noting that a re-emerging superconducting phase between 10.5 GPa and 13.2 GPa was observed experimentallySun et al. (2012). Given the intrinsic phase separation in the superconducting samples, it is possible that the vacancy order survives and the Néel-AM order becomes present under this pressure. In such case, there are at least three possible scenarios relevant to the SC. In the first scenario, the superconducting phase is different from the Néel-AM phase. Then the system becomes natural realization of a superconductor in close proximity of altermagnetism, where finite-momentum Cooper pairing may occurZhang et al. (2024); Chakraborty and Black-Schaffer (2024). It was also proposed that finite-energy pairing in this case may lead to closed superconducting gap and a pair of mirage gapsWei et al. (2024). In the second scenario, the SC emerges from a paramagnetic normal state by suppressing the Néel-AM order; while in the last scenario, the SC emerges directly from a Néel-AM order normal state.
We shall elaborate slightly further on the second and the third cases here. In the second scenario, the time reversal symmetry is present in the normal state, and the pairing is likely to be mediated by residual spin fluctuation from suppressed long range order. Employing random phase approximations (RPA), the calculated spin susceptibility exhibits significant enhancement at zone corners when on-site interactions on Fe atoms are employed, and the relation holds due to the SU(2) symmetry [FIG. 3(e), details to the RPA calculations can be found in SI section I]. Meanwhile, the charge susceptibility is suppressed and featureless. Thus the leading spin fluctuation is antiferromagnetic, favoring a spin-singlet pairing [FIG. 3(f)]. In the third scenario, since the SOC is negligible, the SC gap functions can still be organized by spin characters, giving rise to a singlet channel (SSP) , triplet opposite spin channel (OSP) , and 2 triplet equal spin channels (ESP) and . Since the spatial inversion symmetry is preserved, (,) is satisfied in the Néel-AM phase. Therefore, both ESP channels remain active. In contrast, the spin-splitting close to is order of 100 meV except a few degenerate nodal lines, thus neither SSP nor OSP channel is favored. We have solved the linearized gap equations in this case assuming momentum dependent pairing potential due to spin fluctuation as well. The charge susceptibility remains suppressed, but the transverse spin susceptibility is much more enhanced than the longitudinal spin susceptibility in this case [FIG. 3(b)]. The pairing strengths of the leading SSP or OSP channels are negligibly small compared to those of ESP channels () [FIG. 3(c)], thus a triplet ESP dominates in this case. Previously, a similar case assuming electron-phonon coupling was also studiedLeraand et al. (2025). It was found that electron-phonon coupling can also lead to spin-triplet pairing if the interaction is sufficiently anisotropic.
In conclusion, we have performed systematic first-principles calculations to study the electronic structure and magnetism of Fe4Se5 compounds under high pressure. We argue a possible magnetic-phase transition from the BS-AFM to Néel-AM phase, accompanied by an insulator-metal transition at high pressure before the ordered iron-vacancy structure collapse. The Néel-AM phase is an intrinsic altermagnetism that does not require ligand atoms nor supercell construction to break the inversion or translational symmetry between the spin up/down sublattices. It leads to large spin-splitted band structure as well as spin-polarized Fermi surfaces. Finally, if the re-emerging superconducting phase in KyFe2-xSe2 under high pressure emerges from the Néel-AM normal phase, an ESP triplet pairing is naturally expected.
Acknowledgements.
We would like to thank Lunhui Hu, Hua Chen, Chenchao Xu, Minghu Fang, Hangdong Wang, Yuanfeng Xu, Ming Shi, Yang Liu and Yu Song for the stimulating discussions. This work has been supported by the National Key R&D Program of China (Nos. 2024YFA1408303 & 2022YFA1402202) and the National Natural Science Foundation of China (Nos. 12274364 & 12274109). The calculations were performed on the Quantum Many-Body Computing Cluster at Zhejiang University and High Performance Computing Center of Hangzhou Normal University.References
- Šmejkal et al. (2022a) L. Šmejkal, J. Sinova, and T. Jungwirth, Phys. Rev. X 12, 040501 (2022a).
- Wu and Zhang (2004) C. Wu and S.-C. Zhang, Phys. Rev. Lett. 93, 036403 (2004).
- Wu et al. (2007) C. Wu, K. Sun, E. Fradkin, and S.-C. Zhang, Phys. Rev. B 75, 115103 (2007).
- Šmejkal et al. (2022b) L. Šmejkal, J. Sinova, and T. Jungwirth, Phys. Rev. X 12, 031042 (2022b).
- Chen et al. (2024) X. Chen, J. Ren, Y. Zhu, Y. Yu, A. Zhang, P. Liu, J. Li, Y. Liu, C. Li, and Q. Liu, Phys. Rev. X 14, 031038 (2024).
- Li et al. (2022) J. Li, Q. Yao, L. Wu, Z. Hu, B. Gao, X. Wan, and Q. Liu, Nature Communications 13, 919 (2022).
- Krempaský et al. (2024) J. Krempaský, L. Šmejkal, S. W. D’Souza, M. Hajlaoui, G. Springholz, K. Uhlířová, F. Alarab, P. C. Constantinou, V. Strocov, D. Usanov, W. R. Pudelko, R. González-Hernández, A. Birk Hellenes, Z. Jansa, H. Reichlová, Z. Šobáň, R. D. Gonzalez Betancourt, P. Wadley, J. Sinova, D. Kriegner, J. Minár, J. H. Dil, and T. Jungwirth, Nature 626, 517 (2024).
- Chen et al. (2022) W. Chen, M. Gu, J. Li, P. Wang, and Q. Liu, Phys. Rev. Lett. 129, 276601 (2022).
- Mazin (2022) I. I. Mazin, “Notes on altermagnetism and superconductivity,” (2022), arXiv:2203.05000 [cond-mat.supr-con] .
- Zhang et al. (2024) S.-B. Zhang, L.-H. Hu, and T. Neupert, Nature Communications 15, 1801 (2024).
- Mæland et al. (2024) K. Mæland, B. Brekke, and A. Sudbø, Phys. Rev. B 109, 134515 (2024).
- Wei et al. (2024) M. Wei, L. Xiang, F. Xu, L. Zhang, G. Tang, and J. Wang, Phys. Rev. B 109, L201404 (2024).
- Chakraborty and Black-Schaffer (2024) D. Chakraborty and A. M. Black-Schaffer, Phys. Rev. B 110, L060508 (2024).
- Sukhachov et al. (2025) P. Sukhachov, H. G. Giil, B. Brekke, and J. Linder, Phys. Rev. B 111, L220403 (2025).
- Guo et al. (2010) J. Guo, S. Jin, G. Wang, S. Wang, K. Zhu, T. Zhou, M. He, and X. Chen, Phys. Rev. B 82, 180520 (2010).
- Wang et al. (2011) H.-D. Wang, C.-H. Dong, Z.-J. Li, Q.-H. Mao, S.-S. Zhu, C.-M. Feng, H. Q. Yuan, and M.-H. Fang, Europhysics Letters 93, 47004 (2011).
- Fang et al. (2011) M.-H. Fang, H.-D. Wang, C.-H. Dong, Z.-J. Li, C.-M. Feng, J. Chen, and H. Q. Yuan, Europhysics Letters 94, 27009 (2011).
- Sasmal et al. (2008) K. Sasmal, B. Lv, B. Lorenz, A. M. Guloy, F. Chen, Y.-Y. Xue, and C.-W. Chu, Phys. Rev. Lett. 101, 107007 (2008).
- Bao et al. (2011) W. Bao, Q.-Z. Huang, G.-F. Chen, D.-M. Wang, J.-B. He, and Y.-M. Qiu, Chinese Physics Letters 28, 086104 (2011).
- Cao and Dai (2011) C. Cao and J. Dai, Phys. Rev. Lett. 107, 056401 (2011).
- Yan et al. (2011) X.-W. Yan, M. Gao, Z.-Y. Lu, and T. Xiang, Phys. Rev. B 83, 233205 (2011).
- Ye et al. (2011) F. Ye, S. Chi, W. Bao, X. F. Wang, J. J. Ying, X. H. Chen, H. D. Wang, C. H. Dong, and M. Fang, Phys. Rev. Lett. 107, 137003 (2011).
- Toulemonde et al. (2013) P. Toulemonde, D. Santos-Cottin, C. Lepoittevin, P. Strobel, and J. Marcus, Journal of Physics: Condensed Matter 25, 075703 (2013).
- Bao (2014) W. Bao, Journal of Physics: Condensed Matter 27, 023201 (2014).
- Mangelis et al. (2019) P. Mangelis, R. J. Koch, H. Lei, R. B. Neder, M. T. McDonnell, M. Feygenson, C. Petrovic, A. Lappas, and E. S. Bozin, Phys. Rev. B 100, 094108 (2019).
- Croitori et al. (2020) D. Croitori, I. Filippova, V. Kravtsov, A. Günther, S. Widmann, D. Reuter, H.-A. Krug von Nidda, J. Deisenhofer, A. Loidl, and V. Tsurkan, Phys. Rev. B 101, 054516 (2020).
- Chen et al. (2021) Y. Chen, J. Jiang, H. Yang, P. Dudin, A. Barinov, Z. Liu, H. Wen, L. Yang, and Y. Chen, Nano Research 14, 823 (2021).
- Chen et al. (2014) T.-K. Chen, C.-C. Chang, H.-H. Chang, A.-H. Fang, C.-H. Wang, W.-H. Chao, C.-M. Tseng, Y.-C. Lee, Y.-R. Wu, M.-H. Wen, H.-Y. Tang, F.-R. Chen, M.-J. Wang, M.-K. Wu, and D. V. Dyck, Proceedings of the National Academy of Sciences 111, 63 (2014), https://www.pnas.org/doi/pdf/10.1073/pnas.1321160111 .
- Gao et al. (2017) M. Gao, X. Kong, X.-W. Yan, Z.-Y. Lu, and T. Xiang, Phys. Rev. B 95, 174523 (2017).
- Yeh et al. (2020) K.-Y. Yeh, Y.-R. Chen, T.-S. Lo, P. M. Wu, M.-J. Wang, K.-S. Chang-Liao, and M.-K. Wu, Frontiers in Physics 8 (2020), 10.3389/fphy.2020.567054.
- Sun et al. (2012) L. Sun, X.-J. Chen, J. Guo, P. Gao, Q.-Z. Huang, H. Wang, M. Fang, X. Chen, G. Chen, Q. Wu, C. Zhang, D. Gu, X. Dong, L. Wang, K. Yang, A. Li, X. Dai, H.-k. Mao, and Z. Zhao, Nature 483, 67 (2012).
- Ksenofontov et al. (2012) V. Ksenofontov, S. A. Medvedev, L. M. Schoop, G. Wortmann, T. Palasyuk, V. Tsurkan, J. Deisenhofer, A. Loidl, and C. Felser, Phys. Rev. B 85, 214519 (2012).
- Chi et al. (2013) S. Chi, F. Ye, W. Bao, M. Fang, H. D. Wang, C. H. Dong, A. T. Savici, G. E. Granroth, M. B. Stone, and R. S. Fishman, Phys. Rev. B 87, 100501 (2013).
- Cao et al. (2011) C. Cao, M. Fang, and J. Dai, “Block spin magnetic phase transition of ayfe1.6se2 under high pressure,” (2011), arXiv:1108.4322 [cond-mat.supr-con] .
- Kotliar et al. (2006) G. Kotliar, S. Y. Savrasov, K. Haule, V. S. Oudovenko, O. Parcollet, and C. A. Marianetti, Rev. Mod. Phys. 78, 865 (2006).
- Haule (2007) K. Haule, Phys. Rev. B 75, 155113 (2007).
- Šmejkal et al. (2020) L. Šmejkal, R. González-Hernández, T. Jungwirth, and J. Sinova, Science Advances 6, eaaz8809 (2020).
- Shao et al. (2021) D.-F. Shao, S.-H. Zhang, M. Li, C.-B. Eom, and E. Y. Tsymbal, Nature Communications 12, 7061 (2021).
- Xu et al. (2025) C. Xu, S. Wu, G.-X. Zhi, G. Cao, J. Dai, C. Cao, X. Wang, and H.-Q. Lin, Nature Communications 16, 3114 (2025).
- Ke et al. (2012) L. Ke, M. van Schilfgaarde, and V. Antropov, Phys. Rev. B 86, 020402 (2012).
- Mazin et al. (2008) I. I. Mazin, M. D. Johannes, L. Boeri, K. Koepernik, and D. J. Singh, Phys. Rev. B 78, 085104 (2008).
- Ye et al. (2014) F. Ye, W. Bao, S.-X. Chi, M.-H. Fang, H.-D. Wang, Q.-H. Mao, J.-C. Wang, J.-J. Liu, S. Jie-Ming, A. M. d. Santos, and J. J. Molaison, Chinese Physics Letters 31, 127401 (2014).
- Tsuchiya et al. (2017) Y. Tsuchiya, S. Ikeda, X.-W. Zhang, S. Kishimoto, T. Kikegawa, N. Hirao, S. I. Kawaguchi, Y. Ohishi, and H. Kobayashi, Journal of the Physical Society of Japan 86, 033705 (2017), https://doi.org/10.7566/JPSJ.86.033705 .
- Bendele et al. (2013) M. Bendele, C. Marini, B. Joseph, G. M. Pierantozzi, A. S. Caporale, A. Bianconi, E. Pomjakushina, K. Conder, A. Krzton-Maziopa, T. Irifune, T. Shinmei, S. Pascarelli, P. Dore, N. L. Saini, and P. Postorino, Phys. Rev. B 88, 180506 (2013).
- Yamamoto et al. (2016) Y. Yamamoto, H. Yamaoka, M. Tanaka, H. Okazaki, T. Ozaki, Y. Takano, J.-F. Lin, H. Fujita, T. Kagayama, K. Shimizu, N. Hiraoka, H. Ishii, Y.-F. Liao, K.-D. Tsuei, and J. Mizuki, Scientific Reports 6, 30946 (2016).
- Yu et al. (2011) R. Yu, P. Goswami, and Q. Si, Phys. Rev. B 84, 094451 (2011).
- Jaeschke-Ubiergo et al. (2024) R. Jaeschke-Ubiergo, V. K. Bharadwaj, T. Jungwirth, L. Šmejkal, and J. Sinova, Phys. Rev. B 109, 094425 (2024).
- Svitlyk et al. (2011) V. Svitlyk, D. Chernyshov, E. Pomjakushina, A. Krzton-Maziopa, K. Conder, V. Pomjakushin, and V. Dmitriev, Inorganic Chemistry 50, 10703 (2011), pMID: 21988233, https://doi.org/10.1021/ic201160y .
- Leraand et al. (2025) K. Leraand, K. Mæland, and A. Sudbø, “Phonon-mediated spin-polarized superconductivity in altermagnets,” (2025), arXiv:2502.08704 [cond-mat.supr-con] .